Week of 190909

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Portal
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes



  • local: Miro (DB, Chair), Borja (Monitoring), Vincent (Security), Kate (WLCG), Julia (WLCG), Gavin (Compute), Frank (Atlas)
  • remote: Christoph (CMS), Darren (RAL), Xavier (KIT), Marcelo (CNAF), Mike (ASGC), Onno (NL-T1), Di (TRIUMF), Christian (NDGF), Vincenzo (EGI), Xin (BNL)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities:
      • tail of ongoing reprocessing campaign (using inputs from tape) ~20K files left at KIT

  • CMS reports ( raw view) -
    • Nothing major to report
    • Minor thing: Lost 10 MC files from Tape at JINR GGUS:143034 - Lost statistics already invalidated in CMS bookkeeping.

  • ALICE -
    • NTR, at least until this afternoon

  • LHCb reports ( raw view) -
    • Activity:
      • MC, user jobs and data restripping.
      • Massive staging at all T1
    • Issues:
      • RAL:
        • GGUS:142350; Under investigation. User jobs increased, no queue. Issue seems to continue.

Sites / Services round table:

  • ASGC:
    • Some checksum issue when stage out file yesterday. The bigpanda was in maintenance this morning, and we will keep watching if this issue persists.
  • BNL: NTR
  • EGI: NTR
  • IN2P3: NC
  • KIT:
  • NDGF:
    • During dcache headnodes update last week: dcache door disappeared when it was single point of failure. Draining was immediately stopped, but the failed headnode could not be resurrected without a restart.
    • Pool upgrades at HPC2N tomorrow. Some ATLAS data unavailable for a few minutes.
    • Broken tape found at UCPH. Leader pin had broken off. Files reported missing.
  • NL-T1:
    • Next week (Sept 16), we'll be upgrading dCache from 4.2 to 5.2, to support Xrootd TPC. GOCDB
  • NRC-KI: NC
  • OSG: NC
  • PIC: NC
  • RAL: NTR

  • CERN computing services:
    • OTG:0052156 : Monday 16th September (10.00 - 14.00). ce503, ce504, ce515, ce516 will be deleted (from Wigner) and recreated in Meyrin. They will be drained, but experiments are recommended to stop submitting to these asap.
  • CERN storage services: NTR
  • CERN databases: NTR
  • Monitoring: NTR
  • MW Officer: NTR
  • Networks: NTR
  • Security: NTR


Edit | Attach | Watch | Print version | History: r19 < r18 < r17 < r16 < r15 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r19 - 2019-09-09 - MiroslavPotocky
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback